OpenStudy (anonymous):

Please Help. The cold water faucet can fill a tub in 30 min. With both the cold and hot water faucets running, the tub can be filled in 20 min. How long does it take the hot water faucet alone to fill the tub.

OpenStudy (anonymous):

Let t = time for warm to fill tub Cold faucet fills 1/30 tub a minute. Both faucets fills 1/20 tub a minute. 1/30 + t = 1/20 t = 1/60 Therefore, 60 minutes to fill a tub with only hot water.